OpenStudy (anonymous):

Mel Needed Two Pieces Of Wood . He Needed One Piece That To Be 1 5/6 Feet Long And The Other Piece To Be 2 3/4 Feet Long. He Cut Both Lengths From A Piece Of Wood That Was 10 Feet Long. How Much Wood Does He Have Left.?

OpenStudy (anonymous):

The best way to approach this is by first adding the whole numbers, then adding the fractions with a common denominator. 1+2= 3 so we'll keep that in mind and move on to the fractions. you want 5/6 and 3/4 to have common denominators so you can add them easily. the fastest way to get a common denominator is to multiply each by the others denominator. \[\frac{ 5 }{ 6 }*4/4 = 20/24\] 3/4 * 6/6= 18/24 Note that each time you multiply by a number that is equal to 1 so you aren't actually changing the answer (4/4 = 1 and 6/6 = 1) then just add the two new fractions together: 20/24 + 18/24 = 38/24 and simplify =19/12 There are 12 inches to a foot so 19/12 = 12/12 + 7/12 so its 1 foot and 7 inches. Now add that to the 3 feet from earlier and you get that the total length he cuts is 4 ft 7 inches. But you want how much wood is left so subtract that from 10 10 ft - 4 ft 7 in = 5 ft 5 in